\begin{thebibliography}{10}

\bibitem{rajpurkar2022ai}
Pranav Rajpurkar, Emma Chen, Oishi Banerjee, and Eric~J Topol.
\newblock {AI}in health and medicine.
\newblock {\em Nature medicine}, 28(1):31--38, 2022.

\bibitem{wu2019deep}
Nan Wu, Jason Phang, Jungkyu Park, Yiqiu Shen, Zhe Huang, Masha Zorin, Stanis{\l}aw Jastrz{\k{e}}bski, Thibault F{\'e}vry, Joe Katsnelson, Eric Kim, et~al.
\newblock Deep neural networks improve radiologists’ performance in breast cancer screening.
\newblock {\em IEEE transactions on medical imaging}, 39(4):1184--1194, 2019.

\bibitem{mckinney2020evaluation}
Scott~Mayer McKinney, Marcin Sieniek, Varun Godbole, Jonathan Godwin, Natasha Antropova, Hutan Ashrafian, Trevor Back, Mary Chesus, Greg~S Corrado, Ara Darzi, et~al.
\newblock International evaluation of an {AI}system for breast cancer screening.
\newblock {\em Nature}, 577(7788):89--94, 2020.

\bibitem{cross2024bias}
James~L Cross, Michael~A Choma, and John~A Onofrey.
\newblock Bias in medical ai: Implications for clinical decision-making.
\newblock {\em PLOS Digital Health}, 3(11):e0000651, 2024.

\bibitem{li2021algorithmic}
Lan Li, Tina Lassiter, Joohee Oh, and Min~Kyung Lee.
\newblock Algorithmic hiring in practice: Recruiter and hr professional's perspectives on {AI}use in hiring.
\newblock In {\em Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society}, AIES '21, page 166–176, New York, NY, USA, 2021. Association for Computing Machinery.

\bibitem{lawton2022ai}
G~Lawton.
\newblock {AI}hiring bias: Everything you need to know. techtarget, 2022.

\bibitem{ccj2024ai}
{Council on Criminal Justice}.
\newblock The implications of {AI}for criminal justice: Key takeaways from a convening of leading stakeholders.
\newblock Hosted by the Stanford Criminal Justice Center, Stanford University School of Law, October 2024.

\bibitem{rigano2019ai}
Christopher Rigano.
\newblock Using artificial intelligence to address criminal justice needs.
\newblock {\em National Institute of Justice Journal}, 280(1-10):17, 2019.

\bibitem{zavrsnik2020ai}
Ale{\v{s}} Zavr{\v{s}}nik.
\newblock Criminal justice, artificial intelligence systems, and human rights.
\newblock In {\em ERA forum}, volume 20(4), pages 567--583. Springer, 2020.

\bibitem{sushina2019ai}
Tatyana Sushina and Andrew Sobenin.
\newblock Artificial intelligence in the criminal justice system: leading trends and possibilities.
\newblock In {\em 6th International Conference on Social, economic, and academic leadership (ICSEAL-6-2019)}, pages 432--437. Atlantis Press, 2020.

\bibitem{oecd2019}
{OECD}.
\newblock Recommendation of the council on artificial intelligence.
\newblock \url{https://tinyurl.com/dkkjef5j}, 2019.
\newblock OECD/LEGAL/0449, at 7--8 \& sec. 1.3.

\bibitem{euaiact2021_13_25}
{European Commission}.
\newblock Proposal for a regulation of the european parliament and of the council laying down harmonized rules on artificial intelligence.
\newblock \url{https://tinyurl.com/vmyaxp4d}, 2021.
\newblock COM (2021) 206 final, Apr. 21, 2021, articles 13, 25.

\bibitem{gdpr2016}
{European Parliament and Council}.
\newblock Regulation (eu) 2016/679 on the protection of natural persons with regard to the processing of personal data (gdpr).
\newblock \url{https://tinyurl.com/37z3c5hp}, 2016.
\newblock 2016 O.J. (L 119) 1, 46, art. 22.

\bibitem{baker2023explainable}
Stephanie Baker and Wei Xiang.
\newblock Explainable {AI}is responsible ai: How explainability creates trustworthy and socially responsible artificial intelligence.
\newblock {\em arXiv preprint arXiv:2312.01555}, 2023.

\bibitem{yampolskiy2019}
Roman~V Yampolskiy.
\newblock Unexplainability and incomprehensibility of {AI}.
\newblock {\em Journal of Artificial Intelligence and Consciousness}, 7(02):277--291, 2020.

\bibitem{lipton2018mythos}
Zachary~C Lipton.
\newblock The mythos of model interpretability: In machine learning, the concept of interpretability is both important and slippery.
\newblock {\em Queue}, 16(3):31--57, 2018.

\bibitem{rozen2023case}
Hofit~Wasserman Rozen, Niva Elkin-Koren, and Ran Gilad-Bachrach.
\newblock The case against explainability.
\newblock {\em arXiv preprint arXiv:2305.12167}, 2023.

\bibitem{baker2025monitoring}
Bowen Baker, Joost Huizinga, Leo Gao, Zehao Dou, Melody~Y Guan, Aleksander Madry, Wojciech Zaremba, Jakub Pachocki, and David Farhi.
\newblock Monitoring reasoning models for misbehavior and the risks of promoting obfuscation.
\newblock {\em arXiv preprint arXiv:2503.11926}, 2025.

\bibitem{martens2025beware}
David Martens, Galit Shmueli, Theodoros Evgeniou, Kevin Bauer, Christian Janiesch, Stefan Feuerriegel, Sebastian Gabel, Sofie Goethals, Travis Greene, Nadja Klein, et~al.
\newblock Beware of" explanations" of {AI}.
\newblock {\em arXiv preprint arXiv:2504.06791}, 2025.

\bibitem{cabitza2023painting}
Federico Cabitza, Andrea Campagner, Chiara Natali, Enea Parimbelli, Luca Ronzio, and Matteo Cameli.
\newblock Painting the black box white: experimental findings from applying {XAI} to an {ECG} reading setting.
\newblock {\em Machine Learning and Knowledge Extraction}, 5(1):269--286, 2023.

\bibitem{cabitza2024explanations}
Federico Cabitza, Caterina Fregosi, Andrea Campagner, and Chiara Natali.
\newblock Explanations considered harmful: the impact of misleading explanations on accuracy in hybrid human-{AI} decision making.
\newblock In {\em World conference on explainable artificial intelligence}, pages 255--269. Springer, 2024.

\bibitem{rudin2019stop}
Cynthia Rudin.
\newblock Stop explaining black box machine learning models for high stakes decisions and use interpretable models instead.
\newblock {\em Nature machine intelligence}, 1(5):206--215, 2019.

\bibitem{rozen2024lost}
Hofit Wasserman-Rozen, Ran Gilad-Bachrach, and Niva Elkin-Koren.
\newblock Lost in translation: the limits of explainability in {AI}.
\newblock {\em Cardozo Arts \& Ent. LJ}, 42:391, 2024.

\bibitem{ghassemi2021false}
Marzyeh Ghassemi, Luke Oakden-Rayner, and Andrew~L Beam.
\newblock The false hope of current approaches to explainable artificial intelligence in health care.
\newblock {\em The Lancet Digital Health}, 3(11):e745--e750, 2021.

\bibitem{nicholls2024pentesting}
Mark Nicholls.
\newblock Types of penetration testing: Black box, white box \& grey box, 2024.

\bibitem{nidhra2012testing}
Srinivas Nidhra and Jagruthi Dondeti.
\newblock Black box and white box testing techniques-a literature review.
\newblock {\em International Journal of Embedded Systems and Applications (IJESA)}, 2(2):29--50, 2012.

\bibitem{khan2012comparative}
Mohd~Ehmer Khan and Farmeena Khan.
\newblock A comparative study of white box, black box and grey box testing techniques.
\newblock {\em International Journal of Advanced Computer Science and Applications}, 3(6), 2012.

\bibitem{euaiact2021}
{European Commission}.
\newblock Proposal for a regulation of the european parliament and of the council laying down harmonized rules on artificial intelligence.
\newblock \url{https://tinyurl.com/vmyaxp4d}, 2021.
\newblock COM (2021) 206 final, Apr. 21, 2021.

\bibitem{aaa2022}
Algorithmic accountability act of 2022, h.r. 6580, 117th cong., 2022.

\bibitem{fabris2025fairness}
Alessandro Fabris, Nina Baranowska, Matthew~J Dennis, David Graus, Philipp Hacker, Jorge Saldivar, Frederik Zuiderveen~Borgesius, and Asia~J Biega.
\newblock Fairness and bias in algorithmic hiring: A multidisciplinary survey.
\newblock {\em ACM Transactions on Intelligent Systems and Technology}, 16(1):1--54, 2025.

\bibitem{rigotti2024fairness}
Carlotta Rigotti and Eduard Fosch-Villaronga.
\newblock Fairness, {AI} \& recruitment.
\newblock {\em Computer Law \& Security Review}, 53:105966, 2024.

\bibitem{myllyaho2021validation}
Lalli Myllyaho, Mikko Raatikainen, Tomi M{\"a}nnist{\"o}, Tommi Mikkonen, and Jukka~K Nurminen.
\newblock Systematic literature review of validation methods for {AI} systems.
\newblock {\em arXiv preprint arXiv:2107.12190}, 2021.

\bibitem{fda2024devices}
{U.S. Food \& Drug Admin. (FDA)}.
\newblock {AI} and machine learning {(AI/ML)}-enabled medical devices.
\newblock \url{https://tinyurl.com/24kpf6zy}, 2024.

\bibitem{microsoft2023azure}
{Microsoft}.
\newblock Recommendations for designing for simplicity and efficiency: Azure well-architected framework reliability, 2023.
\newblock December 2, 2023.

\bibitem{dau2025asoe}
{Defense Acquisition University}.
\newblock Affordable system operational effectiveness (asoe) model.
\newblock \url{https://www.dau.edu/}.

\bibitem{federalreserve2011}
{Board of Governors of the Federal Reserve System}.
\newblock Guidance on model risk management, sr 11-7.
\newblock \url{https://tinyurl.com/kcbpyxk4}, April 2011.

\bibitem{cac2022}
{Cyberspace Administration of China}.
\newblock Provisions on the management of algorithmic recommendations in internet information services, 2022.
\newblock Promulgated Jan. 4, 2022; Effective Mar. 1, 2022, arts. 7--8.

\bibitem{cac2023}
{Cyberspace Administration of China}.
\newblock Measures for the management of generative artificial intelligence services (draft for comment), April 2023.
\newblock Art. 17.

\bibitem{fda2021gmlp}
{U.S. Food and Drug Administration}.
\newblock Good machine learning practice for medical device development: Guiding principles.
\newblock \url{https://tinyurl.com/yrnfprct}, October 2021.
\newblock Art. 10.

\bibitem{eastman2023vehicles}
Brittany Eastman, Shay Collins, Ryan Jones, JJ~Martin, Marjory~S Blumenthal, and Karlyn~D Stanley.
\newblock A comparative look at various countries' legal regimes governing automated vehicles.
\newblock {\em JL \& Mobility}, page~1, 2023.

\bibitem{ebert2019validation}
Christof Ebert and Michael Weyrich.
\newblock Validation of automated and autonomous vehicles.
\newblock {\em ATZelectronics worldwide}, 14(9):26--31, 2019.

\bibitem{anderson2016autonomous}
James~M Anderson, Kalra Nidhi, Karlyn~D Stanley, Paul Sorensen, Constantine Samaras, and Oluwatobi~A Oluwatola.
\newblock {\em Autonomous vehicle technology: A guide for policymakers}.
\newblock Rand Corporation, 2016.

\bibitem{hoppenheit2015inventory}
Steffi Hoppenheit and Willibald~A G{\"u}nthner.
\newblock Identifying main drivers on inventory using regression analysis.
\newblock In {\em Operational Excellence in Logistics and Supply Chains: Optimization Methods, Data-driven Approaches and Security Insights. Proceedings of the Hamburg International Conference of Logistics (HICL), Vol. 22}, pages 141--169. Berlin: epubli GmbH, 2015.

\bibitem{messenger2004minoxidil}
AG~Messenger and J~Rundegren.
\newblock Minoxidil: mechanisms of action on hair growth.
\newblock {\em British journal of dermatology}, 150(2):186--194, 2004.

\bibitem{sahin2024blackbox}
Emrullah Şahin, N.N. Arslan, and D.~Özdemir.
\newblock Unlocking the black box: An in-depth review on interpretability, explainability, and reliability in deep learning.
\newblock {\em Neural Computing and Applications}, 37:859--965, 2025.

\bibitem{turbe2023posthoc}
Hugues Turb{\'e}, Mina Bjelogrlic, Christian Lovis, and Gianmarco Mengaldo.
\newblock Evaluation of post-hoc interpretability methods in time-series classification.
\newblock {\em Nature Machine Intelligence}, 5(3):250--260, 2023.

\bibitem{carloni2023causality}
Gianluca Carloni, Andrea Berti, and Sara Colantonio.
\newblock The role of causality in explainable artificial intelligence.
\newblock {\em Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery}, 15(2):e70015, 2025.

\bibitem{kosasih2024supplychain}
Edward~Elson Kosasih, Emmanuel Papadakis, George Baryannis, and Alexandra Brintrup.
\newblock A review of explainable artificial intelligence in supply chain management using neurosymbolic approaches.
\newblock {\em International Journal of Production Research}, 62(4):1510--1540, 2024.

\bibitem{finance2024risk}
Qi~Shen.
\newblock Ai-driven financial risk management systems: Enhancing predictive capabilities and operational efficiency.
\newblock {\em Applied and Computational Engineering}, 69:134--139, 2024.

\bibitem{fassi2024validation}
Sammy Chouffani El~Fassi, Adonis Abdullah, Ying Fang, Sarabesh Natarajan, Awab~Bin Masroor, Naya Kayali, Simran Prakash, and Gail~E Henderson.
\newblock Not all {AI} health tools with regulatory authorization are clinically validated.
\newblock {\em Nature Medicine}, 30(10):2718--2720, 2024.

\bibitem{coe2024framework}
{Council of Europe}.
\newblock Framework convention on artificial intelligence and human rights, democracy and the rule of law, 2024.

\bibitem{uk2023approach}
{U.K. Secretary of State for Science, Innovation \& Technology}.
\newblock A pro-innovation approach to {AI} regulation, March 2023.

\bibitem{execorder14960_2023}
Joseph Biden.
\newblock Executive order no. 14,960 on the safe, secure, and trustworthy development and use of artificial intelligence.
\newblock \url{https://tinyurl.com/49jfr2vz}, November 2023.
\newblock 88 Fed. Reg. 74,899, Secs. 2(a), 4.1(a), 4.2, 4.3.

\bibitem{trump2025eo14179}
Donald~J. Trump.
\newblock Executive order 14179: Removing barriers to american leadership in artificial intelligence.
\newblock \url{https://tinyurl.com/4sxmyfsx/}, January 2025.

\bibitem{omb2025m2521}
{Office of Management and Budget}.
\newblock M-25-21: Accelerating federal use of {AI} through innovation, governance, and public trust.
\newblock \url{https://tinyurl.com/x2w3w6sb}, April 2025.

\bibitem{trump2020trustworthy}
Donald~J. Trump.
\newblock Promoting the use of trustworthy artificial intelligence in the federal government.
\newblock \url{https://tinyurl.com/5fyu3z28}, December 2020.

\bibitem{california2023ab331}
{California}.
\newblock Ab 331 bill, section 22756.3(a)(3), 2023.

\bibitem{nist2024ai6001}
{National Institute of Standards and Technology}.
\newblock {NIST} {AI} 600-1 artificial intelligence risk management framework: Generative artificial intelligence profile, 2024.
\newblock NIST Trustworthy and Responsible AI.

\bibitem{frost2024interpretable}
Nave Frost, Zachary Lipton, Yishay Mansour, and Michal Moshkovitz.
\newblock Partially interpretable models with guarantees on coverage and accuracy.
\newblock In {\em International conference on algorithmic learning theory}, pages 590--613. PMLR, 2024.

\bibitem{gramegna2021shaplime}
Alex Gramegna and Paolo Giudici.
\newblock Shap and lime: an evaluation of discriminative power in credit risk.
\newblock {\em Frontiers in Artificial Intelligence}, 4:752558, 2021.

\bibitem{slack2020fooling}
Dylan Slack, Sophie Hilgard, Emily Jia, Sameer Singh, and Himabindu Lakkaraju.
\newblock Fooling lime and shap: Adversarial attacks on post hoc explanation methods.
\newblock In {\em Proceedings of the AAAI/ACM Conference on AI, Ethics, and Society}, pages 180--186, 2020.

\bibitem{chander2024tai}
Bhanu Chander, Chinju John, Lekha Warrier, and Kumaravelan Gopalakrishnan.
\newblock Toward trustworthy artificial intelligence (tai) in the context of explainability and robustness.
\newblock {\em ACM Computing Surveys}, 57(6):1--49, 2025.

\bibitem{ranchordas2021experimental}
Sofia Ranchord{\'a}s.
\newblock Experimental regulations and regulatory sandboxes: Law without order?
\newblock {\em Law and method}, 2021, 2021.

\bibitem{poursabzi2021interpretable}
Forough Poursabzi-Sangdeh, Daniel~G Goldstein, Jake~M Hofman, Jennifer~Wortman Wortman~Vaughan, and Hanna Wallach.
\newblock Manipulating and measuring model interpretability.
\newblock In {\em Proceedings of the 2021 CHI conference on human factors in computing systems}, pages 1--52, 2021.

\bibitem{kaur2020interpreting}
Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman~Vaughan.
\newblock Interpreting interpretability: understanding data scientists' use of interpretability tools for machine learning.
\newblock In {\em Proceedings of the 2020 CHI conference on human factors in computing systems}, pages 1--14, 2020.

\bibitem{elaziz2021metaheuristic}
Rabia Aziz, CK~Verma, and Namita Srivastava.
\newblock Artificial neural network classification of high dimensional data with novel optimization approach of dimension reduction.
\newblock {\em Annals of Data Science}, 5(4):615--635, 2018.

\bibitem{aziz2018ann}
Rabia Aziz, C.K. Verma, and Namita Srivastava.
\newblock Artificial neural network classification of high-dimensional data with novel optimization approach of dimension reduction.
\newblock {\em Annals of Data Science}, 5:615, 2018.

\bibitem{mikalef2022darkside}
Patrick Mikalef, Kieran Conboy, Jenny~Eriksson Lundstr{\"o}m, and Ale{\v{s}} Popovi{\v{c}}.
\newblock Thinking responsibly about responsible ai and ‘the dark side’of ai.
\newblock {\em European Journal of Information Systems}, 31(3):257--268, 2022.

\bibitem{fuchs2018bias}
Daniel~James Fuchs.
\newblock The dangers of human-like bias in machine-learning algorithms.
\newblock {\em Missouri S\&T’s Peer to Peer}, 2(1):1, 2018.

\bibitem{pasquale2015blackbox}
Frank Pasquale.
\newblock {\em The black box society: The secret algorithms that control money and information}.
\newblock Harvard University Press, 2015.

\bibitem{denzel2024certification}
Philipp Denzel, Stefan Brunner, Yann Billeter, Oliver Forster, Carmen Frischknecht-Gruber, Monika Reif, Frank-Peter Schilling, Joanna Weng, Ricardo Chavarriaga, Amin Amini, et~al.
\newblock Towards the certification of ai-based systems.
\newblock In {\em 2024 11th IEEE Swiss Conference on Data Science (SDS)}, pages 84--91. IEEE, 2024.

\bibitem{europeancommission2020safety}
{European Commission}.
\newblock Report on the safety and liability implications of artificial intelligence, the internet of things and robotics, February 2020.
\newblock COM(2020) 64 final, Section 3.

\bibitem{europeancommission2019liability}
{European Commission, Expert Group on Liability and New Technologies}.
\newblock Report on liability for {AI} and other emerging technologies (new technologies formation), November 2019.
\newblock COM (2019), pp. 27--29.

\bibitem{bertolini2021liability}
Andrea Bertolini and Francesca Episcopo.
\newblock The expert group’s report on liability for artificial intelligence and other emerging digital technologies: a critical assessment.
\newblock {\em European Journal of Risk Regulation}, 12(3):644--659, 2021.

\bibitem{myers2020unreliable}
Paul~D Myers, Kenney Ng, Kristen Severson, Uri Kartoun, Wangzhi Dai, Wei Huang, Frederick~A Anderson, and Collin~M Stultz.
\newblock Identifying unreliable predictions in clinical risk models.
\newblock {\em NPJ digital medicine}, 3(1):8, 2020.

\bibitem{pfau2024trustworthy}
Diana Pfau and Alexander Jung.
\newblock Engineering trustworthy {AI}: A developer guide for empirical risk minimization.
\newblock {\em arXiv preprint arXiv:2410.19361}, 2024.

\bibitem{banerjee2020failures}
Debarag~Narayan Banerjee and Sasanka~Sekhar Chanda.
\newblock {AI} failures: A review of underlying issues.
\newblock {\em arXiv preprint arXiv:2008.04073}, 2020.

\bibitem{dakka2021dataquality}
M.A. Dakka, T.V. Nguyen, J.M.M. Hall, S.M. Diakiw, M.~VerMilyea, R.~Linke, M.~Perugini, and D.~Perugini.
\newblock Automated detection of poor-quality data: case studies in healthcare.
\newblock {\em Scientific Reports}, 11(1):18005, 2021.

\bibitem{kancharla2019speed}
Akshitha Kancharla and Akhil Pannala.
\newblock Factors for accelerating the development speed in systems of artificial intelligence.
\newblock Master's thesis, Department of Software Engineering, 2019.

\bibitem{mokander2024audit}
Jakob Mokander, Justin Curl, and Mihir Kshirsagar.
\newblock A blueprint for auditing generative {AI}.
\newblock {\em arXiv preprint arXiv:2407.05338}, 2024.

\bibitem{sivathapandi2022synthetic}
Praveen Sivathapandi, Debasish Paul, and Akila Selvaraj.
\newblock {AI}-generated synthetic data for stress testing financial systems: A machine learning approach to scenario analysis and risk management.
\newblock {\em Journal of Artificial Intelligence Research}, 2:246--287, Jun. 2022.

\bibitem{ccj2024criminal}
{Council on Criminal Justice}.
\newblock The implications of {AI}for criminal justice: Key takeaways from a convening of leading stakeholders.
\newblock Hosted by the Stanford Criminal Justice Center, Stanford University School of Law, October 2024.

\bibitem{taylor2023algorithm}
Isaac Taylor.
\newblock Justice by algorithm: The limits of {AI} in criminal sentencing.
\newblock {\em Criminal Justice Ethics}, 42(3):193--213, 2023.

\bibitem{carlson2017transparency}
Alyssa~M Carlson.
\newblock The need for transparency in the age of predictive sentencing algorithms.
\newblock {\em Iowa L. Rev.}, 103:303, 2017.

\bibitem{jacobsen2023smart}
Gilson Jacobsen and Bruno de~Macedo~Dias.
\newblock Smart dispute resolution: Artificial intelligence to reduce litigation.
\newblock {\em Suprema--Revista de Estudos Constitucionais}, 3(1):391--414, 2023.

\bibitem{soyer2022liability}
Baris Soyer and Andrew Tettenborn.
\newblock Artificial intelligence and civil liability—do we need a new regime?
\newblock {\em International Journal of Law and Information Technology}, 30(4):385--397, 2022.

\bibitem{holweg2022reputation}
Matthias Holweg, Rupert Younger, and Yuni Wen.
\newblock The reputational risks of {AI}.
\newblock {\em California management review insights}, 2022.

\bibitem{stiglitz2000econinfo}
Joseph~E Stiglitz.
\newblock The contributions of the economics of information to twentieth century economics.
\newblock {\em The quarterly journal of economics}, 115(4):1441--1478, 2000.

\bibitem{baek2024researchagent}
Jinheon Baek, Sujay~Kumar Jauhar, Silviu Cucerzan, and Sung~Ju Hwang.
\newblock Researchagent: Iterative research idea generation over scientific literature with large language models.
\newblock {\em arXiv preprint arXiv:2404.07738}, 2024.

\bibitem{denny2019allofus}
Joshua~C Denny, Stephanie~A Devaney, and Kelly~A Gebo.
\newblock The "all of us" research program. reply.
\newblock {\em The New England journal of medicine}, 381(19):1884—1885, November 2019.

\bibitem{iso2022ai23053}
{International Organization for Standardization and International Electrotechnical Commission}.
\newblock Iso/iec 23053:2022 — framework for artificial intelligence ({AI}) systems using machine learning ({ML}), 2022.

\bibitem{oecd2024ai}
{OECD}.
\newblock Policies, data and analysis for trustworthy artificial intelligence.
\newblock \url{https://oecd.ai/en/}, 2024.

\end{thebibliography}

\documentclass[11pt]{article}

\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{authblk}
\usepackage[margin=1in]{geometry}
\usepackage{cite}
\usepackage{xurl}
\usepackage{multicol}

\title{\textbf{Beyond Explainability: The Case for AI Validation}}

\author{
\begin{multicols}{2}
\begin{tabular}{c c}
\begin{tabular}{c}
\textbf{Dalit Ken-Dror Feldman} \\
University of Haifa \\
\texttt{dalitkd@law.haifa.ac.il}
\end{tabular}
\begin{tabular}{c}
\textbf{Daniel Benoliel} \\
University of Haifa \\
\texttt{dbenolie@law.haifa.ac.il}
\end{tabular}
\end{tabular}
\end{multicols}
}

\date{May 2025}

\begin{document}
\maketitle
\begin{multicols}{2}
\begin{abstract}
Artificial Knowledge (AK) systems are transforming decision-making across critical domains such as healthcare, finance, and criminal justice. However, their growing opacity presents governance challenges that current regulatory approaches, focused predominantly on explainability, fail to address adequately. This article argues for a shift toward validation as a central regulatory pillar. Validation, ensuring the reliability, consistency, and robustness of AI outputs, offers a more practical, scalable, and risk-sensitive alternative to explainability, particularly in high-stakes contexts where interpretability may be technically or economically unfeasible. We introduce a typology based on two axes, validity and explainability, classifying AK systems into four categories and exposing the trade-offs between interpretability and output reliability. Drawing on comparative analysis of regulatory approaches in the EU, US, UK, and China, we show how validation can enhance societal trust, fairness, and safety even where explainability is limited. We propose a forward-looking policy framework centered on pre- and post-deployment validation, third-party auditing, harmonized standards, and liability incentives. This framework balances innovation with accountability and provides a governance roadmap for responsibly integrating opaque, high-performing AK systems into society.
\end{abstract}

\noindent \textbf{Keywords:} Artificial Intelligence governance, validation, explainability, algorithmic accountability, AI regulation and policy, trustworthy AI.

\section{Introduction}
The rise of Artificial Knowledge (AK) challenges traditional models of trust, accountability, and governance due to its opaque nature. While AK systems can process vast datasets and generate powerful insights their complexity often prevents full understanding, introducing epistemic uncertainties~\cite{rajpurkar2022ai,wu2019deep,mckinney2020evaluation}.  This becomes a critical concern in fields like healthcare~\cite{cross2024bias}, finance~\cite{li2021algorithmic,lawton2022ai}, and criminal justice~\cite{ccj2024ai,rigano2019ai,zavrsnik2020ai,sushina2019ai}, where unverified outputs impact fundamental rights. This paper advocates shifting the focus from explainability alone to a combination of explainability and validation of AK systems, offering a more practical and scalable approach for integrating AK into high-stakes domains while preserving trustworthiness and safety.

\section{The Problem with Explainability}
The primary focus of this article is one of the most critical challenges facing AK: its inherent opacity~\cite{oecd2019,euaiact2021_13_25,gdpr2016}.  Artificial systems mitigate this issue by embedding explainability into their outputs, enabling trust. Explainability is grounded in the notion that it promotes trustworthy AI applications~\cite{baker2023explainable}.  However, balancing explainability with technical accuracy presents challenges, as simplifying complex models may compromise precision, raising concerns about fairness and accountability~\cite{yampolskiy2019}. Explainable Artificial Intelligence (XAI) aims to enhance the transparency and trustworthiness of complex "black-box" models by offering human-interpretable explanations for their outputs. This is particularly critical in high-stakes domains where AI-driven decisions require both accountability and user trust. However, recent literature reveals fundamental problems: the notion of interpretability is often ambiguous~\cite{lipton2018mythos,rozen2023case},  and many explanation methods are unreliable or misleading~\cite{baker2025monitoring}.  Poor-quality explanations may create only an illusion of understanding, leading to incorrect human decisions, overconfidence in flawed models, or even enabling manipulation~\cite{martens2025beware}.  
Explanations can obscure model behavior, as seen in the "white box paradox", where they increase trust despite incorrect recommendations. Efforts to clarify black-box models may falsely create transparency, undermining usability and introducing errors~\cite{cabitza2023painting}.   Some take it to the next level and call it the 'XAI halo effect'. This occurs when misleading explanations lead users to trust advice without verifying its accuracy~\cite{cabitza2024explanations}.  Thus, misleading explanations can diminish the benefits of transparency. Moreover, relying on explanations rather than designing inherently interpretable models may perpetuate harmful practices~\cite{rudin2019stop,lipton2018mythos}.

\section{The Advantages of Validation}
In the absence of full explainability tools~\cite{rozen2023case,rozen2024lost,ghassemi2021false}, validated data can help close the regulatory gap, with the system's overall utility could compensate for the limitations inherent in its lack of explainability~\cite{nicholls2024pentesting,nidhra2012testing,khan2012comparative}.  
Validation forms the objective foundation for fairness initiatives by ensuring consistent and reliable system outputs. Both the EU AI Act (article 10)\cite{euaiact2021} and the U.S. Algorithmic Accountability Act of 2022\cite{aaa2022} emphasize fairness, which in contexts like hiring platforms under the GDPR, requires validated processes to prevent discriminatory automated decisions~\cite{fabris2025fairness,rigotti2024fairness,gdpr2016}. Validation focuses on outcomes rather than processes, making it a more feasible and scalable solution for ensuring trust in complex AI systems~\cite{myllyaho2021validation,fda2024devices}. Validation outweighs efficiency. Prioritizing efficiency over validity can result in significant risks, if a system is not adequately validated~\cite{microsoft2023azure,dau2025asoe}. 
Validation ensures efficiency does not compromise reliability. For instance, the EU's AI Act (article 16)~\cite{euaiact2021} and the U.S. Federal Reserve’s Guidance on Model Risk Management~\cite{federalreserve2011} both mandate testing and validation of high-risk AI systems to ensure robustness, prioritizing long-term stability over short-term efficiency gains. By requiring validation, policymakers establish clear benchmarks for assessing AK reliability. For example, China’s AI regulations require validation reports for high-risk systems, giving regulators evidence to assess compliance (articles 7-8, 2022)(article 17, 2023)~\cite{cac2022,cac2023}.  
Moreover, validation inherently ensures safety, especially in high-stakes areas like healthcare and public safety. The FDA's Good Machine Learning Practices (GMLP)~\cite{fda2021gmlp}, and the EU's AI Act (article 10) mandate rigorous testing to ensure systems operate reliably under diverse conditions. For example, autonomous vehicles undergo simulations and real-world tests, where safety is integral to the validation process, not a separate concern~\cite{eastman2023vehicles,ebert2019validation,anderson2016autonomous}. 
Lastly, validation is crucial for ensuring AI systems perform as intended, offering an objective standard for reliability while supporting values like fairness, accountability, and safety. Frameworks such as the EU's AI Act, the FDA's GMLP, and China's AI regulation emphasize validation as central to AI governance, addressing risks, fostering trust, and ensuring compliance with regulatory and societal priorities.

\section{Discussion: The Validity-Explainability Matrix}

Balancing validity and explainability in AI systems is challenging. Highly explainable systems that lack validity may gain trust while producing unreliable outputs, while valid systems with low explainability can erode trust and accountability. This section explores the trade-offs between validation and explainability. It critiques the current focus on explainability, emphasizing that validity and explainability should be complementary. 
Validation ensures output reliability and consistency, while explainability makes outputs interpretable. Valid-explainable AK achieves reliability consistency and interpretability, but is often limited to simpler models or low-stakes contexts where explainability costs are manageable, such as linear regression models in inventory management~\cite{hoppenheit2015inventory}. 
In contrast, valid-opaque AK is reliable and consistent but lack explainability~\cite{messenger2004minoxidil}. This category often includes deep learning systems, where outputs are generated through non-linear and probabilistic processes~\cite{sahin2024blackbox}.  While these AI systems may achieve high levels of performance, the costs of making them explainabl, through post hoc interpretability tools~\cite{turbe2023posthoc}, documentation, or causal mapping~\cite{carloni2023causality}, can be excessive. Achieving explainability of valid-opaque AK may reduce model’s accuracy, delay deployment, or impose significant technical and financial burdens. This trade-off is particularly pronounced in domains where performance optimization is paramount, such as financial risk modeling or AI-driven supply chains~\cite{kosasih2024supplychain,finance2024risk}. 
Non-valid opaque AK exacerbates these issues. It lacks reliability, consistency, and explainability, imposing the highest risks and costs on society. The opacity of such AI systems makes it difficult to detect errors, while their invalid outputs can propagate cascading failures. For instance, a flawed AI system used in credit scoring may penalize certain demographic groups based on incomplete or biased data. Without explainability, stakeholders cannot identify or challenge the source of these biases, leaving societal harm unaddressed~\cite{ghassemi2021false}. 
Given these dynamics, validation is a more practical and cost-effective approach when explainability is not feasible. It ensures consistent, accurate outputs, even if the internal processes are unclear. In high-stakes areas like healthcare or finance, reliable results are crucial for minimizing risks. For instance, a validated medical diagnostic tool provides trustworthy results despite lacking explainability~\cite{fassi2024validation,ghassemi2021false}.  
Balancing validity and explainability in AI systems presents significant challenges. Highly explainable systems that lack validity may gain trust but produce unreliable outputs, while valid but opaque systems can undermine trust and accountability. The EU, U.S., and China have different approaches to balancing these factors, but validation remains crucial, especially in high-risk areas like healthcare and public safety (EU's AI Act, article 13(3)(b)(ii))~\cite{euaiact2021}. 
It is worth noting that the EU’s Framework Convention on AI of 2024~\cite{coe2024framework}, does not explicitly reference the terms “explainability” or “validity”, perhaps reflecting a growing skepticism about the feasibility of fully explaining AI systems. Nevertheless, the document emphasizes the principles of oversight and transparency (article 8)~\cite{coe2024framework} accountability, responsibility (article 9)~\cite{coe2024framework} and reliability (article 12)~\cite{coe2024framework}. It highlights the importance of maintaining human control over AI processes, even in the absence of a detailed understanding of their mechanics. Comparably, in the UK White Paper, the concept of explainability is mentioned, but the requirement for validity is not addressed~\cite{uk2023approach}. This arguably leaves room for greater emphasis on validation as the more practical alternative in specific contexts.
In contrast, U.S. policy places greater emphasis on validity. In 2023, President Biden issued the Executive Order on the Safe, Secure, and Trustworthy Development and Use of AI, mandating robust evaluation measures to ensure AI systems are reliable, secure, and function as intended (section 2(a))~\cite{execorder14960_2023}. This comprehensive framework governs the development, testing, and deployment of AI technologies (section 4.1(a))~\cite{execorder14960_2023}. The Executive Order emphasizes the importance of validity in AI systems, particularly through its focus on safety, security, and trustworthiness (sections 4.2 and 4.3)~\cite{execorder14960_2023}. Although this Executive Order was canceled by President Trump in 2025~\cite{trump2025eo14179}, the notion of validity remained. The Office of Management and Budget (2025)~\cite{omb2025m2521} highlights the importance of continuous validation of AI systems. Similarly, the 2020 executive order by President Trump~\cite{trump2020trustworthy}, emphasizes the need for AI systems to be "accurate, reliable, and effective" in their deployment, further reinforcing the need for validation in AI applications.
Both validity and explainability are included in the proposed California legislation~\cite{california2023ab331}, as well as in the National Institute of Standards and Technology (NIST) Principles~\cite{nist2024ai6001}, reflecting a broader emphasis on these values for ensuring the reliability, responsibility and trustworthiness of AI systems. The emphasis on validation, though, must not become a justification for perpetuating opacity~\cite{myllyaho2021validation}. Developers should be required to invest in explainability where its costs are reasonable and its benefits, such as fairness, accountability, and human oversight, are significant~\cite{fabris2025fairness,rigotti2024fairness}.  
Policymakers must balance the incentives for validation and explainability through targeted regulation and economic mechanisms. Pre-deployment validation mandates, such as those in the EU AI Act, provide a baseline standard of reliability. Liability frameworks, such as strict liability for system failures, ensure that developers internalize the costs of harm, incentivizing robust testing even when explainability remains elusive. 
AK can be classified along explainability and validity axes, highlighting the trade-offs and challenges in its governance. Table 1 below provides a structured framework for this analysis.

\begin{table*}[t]
\centering
\begin{tabular}{|c|c|c|}
\hline
\textbf{} & \textbf{Explainable} & \textbf{Opaque} \\
\hline
\textbf{Valid} & Valid-Explainable & Valid-Opaque \\
\hline
\textbf{Non-Valid} & Pre-Valid Explainable & Non-Valid Opaque \\
\hline
\end{tabular}
\caption{Typology of Artificial Knowledge Creation}
\label{tab:ak_typology}
\end{table*}

This typology underscores the balance required between explainability and validity, offering a structured approach for governance and regulatory strategies to maximize AK’s benefits while minimizing risks.

\subsection{Valid-Explainable: Reliable, Consistent, Interpretable, and Transparent}
Balancing \textit{validation} and \textit{explainability} in AI presents challenges. While explainable AI can build trust, its complexity can increase costs, and it may not always improve reliability. Validation, however, ensures AI systems perform as intended, making it crucial in high-stakes domains where accuracy is vital. Regulatory frameworks like the EU AI Act, U.S. executive orders, and China’s AI regulations all emphasize validation but differ in their treatment of explainability. Validation, often more cost-effective than explainability, reduces risks, inefficiencies, and societal harm, which can occur from unreliable systems.
In sectors like healthcare, validated systems are more likely to be trusted and adopted. However, this focus on validation shouldn’t justify ignoring explainability where it's feasible, as explainability ensures fairness, accountability, and oversight. Policymakers must balance validation and explainability through targeted regulations, liability frameworks, and market incentives like certification programs, which encourage innovation while maintaining reliability. Ultimately, in some cases, validation may be the more practical approach, fostering trust and safeguarding societal interests when explainability is economically unfeasible. Recent research explores hybrid models that balance both aspects, aiming for high accuracy while maintaining interpretability. For instance, partially interpretable models provide explanations for decisions while ensuring robust performance~\cite{frost2024interpretable,gramegna2021shaplime,slack2020fooling}.   

\subsection{Pre-Valid Explainable: Interpretable but Provisional}
The \textit{Pre-Valid Explainable} quadrant represents AK that prioritizes explainability but has not yet achieved full validation. This knowledge is transparent, interpretable and accessible, allowing stakeholders to understand AI outputs and reasoning, but their findings remain provisional. That is due to limited testing, incomplete training, or insufficient contextual robustness~\cite{chander2024tai}. While unsuitable for high-stakes applications, these systems may be indispensable in exploration, developmental, and speculative contexts.
This approach is particularly fitting for prevalid explainable systems. While interpretable and transparent, these systems are provisional and prone to errors. Traditional regulatory models may prematurely dismiss or overburden them. Experimentalist governance offers a flexible pathway for refinement, with regulatory sandboxes providing a controlled environment to test prevalid systems with reduced compliance requirements~\cite{ranchordas2021experimental}. These sandboxes allow regulators and developers to evaluate system performance, identify risks, and implement improvements without the pressure of full-scale deployment~\cite{ranchordas2021experimental}. 

\subsection{Valid-Opaque: Reliable and Consistent but Inscrutable}
The \textit{Valid-Opaque} quadrant characterizes AK, demonstrating high reliability and consistency but a lack of explainability. These systems consistently produce outputs that meet rigorous standards of accuracy and robustness, yet their internal decision-making processes remain opaque\cite{poursabzi2021interpretable,kaur2020interpreting}. This combination of validity and opacity presents significant challenges in trust, accountability, and governance\cite{kaur2020interpreting,poursabzi2021interpretable}. That is, especially in high-stakes applications where decisions have profound societal implications\cite{kaur2020interpreting, poursabzi2021interpretable,ghassemi2021false}. 
Opacity in valid-opaque systems stem from complex architectures like neural networks and ensemble methods, which are difficult to interpret despite their effectiveness\cite{elaziz2021metaheuristic,aziz2018ann}. This lack of explainability creates accountability challenges, making it hard to identify the sources of errors or biases\cite{mikalef2022darkside,fuchs2018bias}. Pasquale calls this the "accountability gap", where stakeholders can't discern whether issues arise from data, algorithms, or external factors\cite{pasquale2015blackbox}.  
However, balancing validity and opacity requires careful governance. Policymakers must navigate this tension through accountability frameworks, transparency incentives,\cite{denzel2024certification} and liability regulations\cite{europeancommission2020safety, europeancommission2019liability,bertolini2021liability} to ensure responsible and ethical deployment of these systems\cite{europeancommission2020safety}, prioritizing trust, fairness, and accountability\cite{myers2020unreliable}.

\subsection{Non-Valid Opaque: Unreliable and High-Risk}
The \textit{Non-Valid Opaque} quadrant represents the most concerning and problematic category of AK: artificial intelligence systems that produce unreliable outputs while remaining inscrutable.  These outputs fail to meet empirical standards of accuracy, consistency, and robustness, and their opacity prevents stakeholders from diagnosing or addressing underlying flaws\cite{pfau2024trustworthy,banerjee2020failures}. This combination of unreliability and opacity exacerbates the externalities of AK, undermining trust, equity, and the efficiency of broader knowledge systems. 
Non-valid opaque systems often arise from governance failures, such as poor-quality training data, inadequate validation, and misaligned incentives\cite{dakka2021dataquality}. These systems, developed under market pressures or weak oversight\cite{kancharla2019speed}, can perpetuate biases and errors. Unlike valid-opaque systems, which can be mitigated with interpretability tools, non-valid opaque systems require proactive controls, such as pre-deployment audits and stress testing, to prevent harm\cite{mokander2024audit,sivathapandi2022synthetic}. In some cases, outright prohibition may be necessary to protect public safety and equity. These systems also erode public trust in AI, creating long-term societal costs, particularly for marginalized groups\cite{ccj2024criminal,taylor2023algorithm, carlson2017transparency}. 
The economic impact is also significant, as non-valid opaque systems lead to litigation\cite{jacobsen2023smart}, reputational damage\cite{soyer2022liability,holweg2022reputation}, and hinder innovation\cite{stiglitz2000econinfo}. They prevent iterative improvements, reinforcing a cycle of unreliability\cite{baek2024researchagent}. Thus, addressing non-valid opaque systems requires robust regulation, accountability frameworks, and, in some cases, prohibition, to ensure that AI development aligns with societal values of trust, equity, and innovation.

\section{Conclusions and Future Research}
In conclusion, integrating AK systems into high-stakes domains introduces significant trust, accountability, and governance challenges. While explainability is essential, the focus should also shift toward validating the reliability and consistency of these systems' outputs. Relying on explanations rather than designing inherently interpretable models may perpetuate harmful practices; however, validation, which ensures consistency and accuracy over time, can address these issues by providing a more reliable foundation for trust and decision-making. This approach offers a practical pathway for adopting AK in healthcare, finance, and criminal justice without undermining societal values. 
To effectively govern AK systems, a comprehensive, risk-based regulatory framework is required to integrate legal, economic, and institutional mechanisms. Such a framework should mandate both pre-deployment and post-deployment validation. Such a framework would establish independent third-party bodies for the impartial evaluation of high-risk systems and support the development of publicly funded validation infrastructure. The latter include standardized datasets, data repositories, and computational resources, particularly to ensure access for small and medium enterprises\cite{denny2019allofus}.   
Importantly, validation should encompass functional, empirical, and normative dimensions, thereby ensuring technical robustness, fairness, and legal compliance. These mechanisms should be complemented by fallback procedures, such as human oversight in sensitive applications, and transparency requirements, including documentation of data sources, methodologies, and system limitations. Harmonization of validation standards through international bodies such as ISO and the OECD\cite{iso2022ai23053, oecd2024ai}, alongside the standardization of fairness metrics and tools, will further reduce compliance burdens and enhance global trust. Additional measures, liability regimes, certification schemes, public-private partnerships for fairness audits, and mandatory fairness and ethical impact assessments, can further align innovation with accountability. Together, these measures create a comprehensive policy framework that aligns innovation with accountability, ensuring that AK systems are reliable, fair, and trustworthy across diverse contexts.
Future research should explore methods for adequate validation of AK systems, particularly in complex, high-risk contexts. Additionally, developing hybrid models that balance explainability and performance and address AK's governance implications will be critical for ensuring these systems' responsible and ethical deployment. Further work is required to develop and refine the regulatory and policy recommendations outlined above, establishing a robust governance framework for AK systems that balances innovation with accountability, safeguards public interests, and fosters societal trust. Similarly, it is essential to explore the limitations of validation in dynamic systems and understand to what extent validity can truly be assessed.

\section*{Author’s Note}
This article is based on the theoretical framework first introduced in Daniel Benoliel, Dalit Ken-Dror Feldman. The social cost of artificial knowledge. University of Illinois Journal of Law, Technology \& Policy, 2025. Forthcoming.

\bibliographystyle{unsrt}
\bibliography{references}
\end{multicols}
\end{document}